This file provides a chronological history of the WFPC2, including:
|Title||Last Updated||Updated by|
|1. Decontaminations||17 April 2008||M. McMaster|
|2. Darks and Superdarks||14 April 2008||M. McMaster|
|3. Focus Changes||5 Sep 2001||S. Baggett, M. Wiggs|
|4. Temperature Changes||5 Sep 2001||S. Baggett, M. Wiggs|
|5. Pipeline Changes||5 Sep 2001||S. Baggett, M. Wiggs|
|6. Safings||6 May 2003||I. Heyer|
|7. Plate Scale Changes||5 Sep 2001||S. Baggett, M. Wiggs|
|8. List of WFPC2 Observations||5 Sep 2001||S. Baggett, M. Wiggs|
|9. Miscellaneous Items
(incl. WFPC2-rel. STSDAS changes)
|5 Sep 2001||S. Baggett, M. Wiggs|
|10. OTFC Keyword Exception Table||5 Sep 2001||S. Baggett, M. Wiggs|
|11. WFPC2 History in Table Format||5 Sep 2001||S. Baggett, M. Wiggs|
It is designed to provide observers with information about the state of the WFPC2 when their observations were taken. The most important change occurred on APRIL 23, 1994, when the operating temperature was changed from -76° C to -88° C in order to minimize the CTE (Charge Transfer Efficiency) problem and the growth of hot pixels. Care should be taken not to mix reference files from one side of the temperature change with data from the other side.
Each item includes a short description of how this might affect the observer. More detailed information is available in:
- WFPC2 Instrument Handbook.
- Links to Documentation, including status reports, workshop proceedings, and more.
- WFPC2 Instrument Science Reports
- WFPC2 Reference Files
- HST Data Handbook
- WFPC2 Decontamination Dates
This history file contains sections sorted by item type, then a table sorted by date.
The WFPC2 has been warmed approximately once a month (every 49 days since Cycle 11) to about +20° C for 6 - 12 hours to remove any buildup of contaminants. This appears to:
1) Restore essentially the full UV sensitivity of the instrument. Recent photometric monitoring results are kept updated in the WWW memo and associated figure. 2) Repair many of the "hot pixels" (see Figure 2 of the Holtzman et al. article).
The calibration programs include regular photometric monitoring of both a standard star (usually GRW+70D5824; alternate is Feige 110) and a standard cluster (Omega Cen = NGC 5139; alternate is NGC 6752), usually done before and after decontamination procedures. Darks are taken at the rate of at least 5 per week (and 5 more before and after decons). More details on the calibration proposals are available in the WFPC2 Instrument Handbook.
From the observer's standpoint the decon dates may be useful in order to:
Make small adjustments to the photometric zeropoints in the UV. Enhancements to SYNPHOT have been made to aid the user in taking these effects into account. Please refer to the ISR98-03: WFPC2 Longterm Photometric Stability and the SYNPHOT section of the WFPC2 Photometry page.
The following link provides a list of the decontamination dates:
DATES OF DECONTAMINATION PROCEDURES.
Typically, OTFR will have used the best available pipeline dark for calibration; OTFR can be accessed via the HST Archive or STARVIEW. To assess the pedigree of the dark, see the WFPC2 Reference Files.
Pipeline darks are normally available for each 5-10 day period. Prior to Aug. 01, 1996, the dark reference files were created from sets of 10 dark frames, 1800 sec. each, (taken over ~2 weeks time). After to Aug. 01, 1996, the darks have been generated by adding hotpixels (as identified from ~5 dark frames, 1800 sec. each, taken within 2 days) to a superdark generated from more than 100 darkframes. Please refer to the Pipeline Section in this document, or to our online Summary of Darks for more details, or to the HISTORY keywords in the headers of the dark reference files themselves.
To improve upon the pipeline dark calibration, one option is to recalibrate manually using a custom dark reference file (generated from, e.g., the shorter 1000 sec. darks that are taken daily and archived only - not used in pipeline). Scripts and cookbook examples of dark generation are available.
Another option is to use a cosmicray rejection routine that can identify the hot pixels (e.g., COSMICRAYS in IRAF), but care should be taken to set the parameters conservatively or you may also clip out the peaks of bright stars. If the data has been dithered, the Drizzle software may be used to remove both cosmic rays and hot pixels.
NOTE: At T = -76° C (i.e., before April 23) the mean dark count was about 2 DN per 1000 seconds. At T = -88° C (i.e., since April 23) the dark count has been about 0.5 DN per 1000 seconds. For an analysis of the dark current versus time, see ISR 01-05.
There are lists of hotpixels available via WWW. Note that there are several flavors: those for use with the STSDAS task "warmpix", plain hot pixel lists, and HDF hotpixel lists.These may also be retrieved in the form of compressed ASCII files via anonymous ftp:
ftp stsci.edu (username) anonymous (password)
NOTE: As an additional aid in hot pixel identification, as of July 1997, more frequent dark frames are being taken. The new Supplemental Darks program (7621, 7712, 7713) is designed to provide more frequent, short (1000s) darks to be used primarily for the identification of hot pixels. Shorter darks are used so that observations can fit into almost any occultation period, making automatic scheduling feasible. The program is designed to complement the Standard Darks program (7620), whose longer individual observations are better suited to produce high-quality pipeline darks and superdarks. This program has become possible after the Second Servicing Mission thanks to the larger storage capacity of the Solid State Recorder. Currently, about 15 darks are taken each week as part of this program. The darks will not be used in standard STScI products, but are available to observers via the Hubble Data Archive.
We also have in the archive superdarks generated from 100 individual dark frames taken from April 23 to October 01, 1994 (NOTE: standard pipeline darks are also superdarks as of Aug. 1996; see previous section for discussion). These 1994 superdarks were not used in the pipeline but may be useful for recalibrating your data in cases where the total exposure time is relatively long (e.g., > 30 minutes) and the background noise is a limiting factor in your science (e.g., faint object detection).
DELIVERED 1994 November (for data taken after April 23 decon)
NOTE 1: the WFPC2 IDT has delivered alternate superdarks, superbiases, and deltadarks. The IDT Reference Files memo describes how these files were generated, the names for retrieval from starview, and how to use them.
NOTE 2: The Dec. 1995 Hubble Deep Field images were processed by the HDF team using the standard HST pipeline (the STSDAS task calwp2) but with different reference files (similar to the IDT superdarks, superbiases, and deltadarks) and a slightly different treatment of the darktime. These alternate reference files are also available from the archive; the HDF Pipeline Reduction page provides some details as well as the filenames of these products. The history records in the reference file headers provide more details as well.
In general, the focus changes have been very minor and are unlikely to have an effect on your science; a short summary table follows. For more details, please refer to the online HST Focus Pages.
2000 Jun. 15 Secondary mirror moved +3.6 microns. (Day 167 19:19) 2000 Jan. 09 Secondary mirror moved +4.2 microns. 1999 Sep. 15 Secondary mirror moved +3 microns. 1998 Jun. 28 Secondary mirror moved -15.2 microns back at end of campaign. 1998 Jun. 04 Secondary mirror moved +16.6 microns for second NIC3 campaign. 1998 Feb. 01 Secondary mirror moved -18.6, at end of NIC3 campaign. 1998 Jan. 12 Secondary mirror moved +21, for NIC3 campaign. 1997 Mar. 18 Secondary mirror moved towards primary by 2.4 microns (077/22:51). 1996 Oct. 30 Secondary moved outwards by 5 microns (day 304, 1700 UT). The average focus position is now ~+2.5 from nominal, compared to -2.5 prior to the move. 1996 Mar. 14 Secondary moved +6 microns to correct for desorption (day 074/1848Z). 1995 Aug. 28 Secondary moved by ~6.5 microns to compensate for OTA desorption. Before the move, WFPC2 ~ 3.5 microns inside optimal focus; move put WFPC2 about 3 microns outside "best focus." Most observers should be unaffected by move. 1995 Feb. 23 Possible anomaly; if your data was taken between - Mar. 03 these dates, and the scientific results might be sensitive to the PSF shape, please contact us. See Note 1, below. 1995 Jan. 15 5 micron change to adjust for desorption. 1994 Jun. 29 5 micron change to adjust for desorption. 1994 Mar. 04 Minor change. 1993 Dec. 28 Limiting factors: Breathing worst case ~ 5-10 microns during a single orbit. Desorption ~ 1 micron/month during cycle 4.
NOTE 1: Faint Object Camera observations taken on Day 58 (Feb. 27) appear blurred. A possible cause of the blurring is a defocus and/or decenter of the secondary mirror, potentially related to the unusual thermal environment of the telescope - which had been pointed very close to the anti-sun direction for several days (ie., most of the telescope structure is in the shadow caused by the wider instrument compartment at the "bottom" of the telescope). WFPC2 observations taken shortly before and during the blurred FOC observations have been inspected for degradation of the PSF. These observations contain very few well-exposed stars in the PC, and therefore our assessment of the PSF properties in WFPC2 is necessarily limited. However, it appears that any degradation in the light distribution is very minor, and the FWHM of the observed stars is close to normal - the measured increase of less than 10% (to 1.8-1.9 PC pixels in F814W) is within the errors of the measurement. The PSF monitoring carried out on Mar. 04 shows nominal results.
However, the _shape_ of the PSF in the WFPC2 observations may be consistent with a defocusing of the secondary mirror of order of 10 micron, the value suggested by the FOC data. Such a defocusing could affect photometry based on small apertures or PSF fitting at the 5-10% level; scientific programs that depend on accurate knowledge of the PSF itself, such as observations of faint objects in the vicinity of bright point sources, would also need to take the possible defocusing into account. On the other hand, morphology and surface photometry of extended sources would be largely unaffected. Information about this event is detailed in the March 1995 STAN.
More general information about the OTA and focus changes may be obtained from the Observatory Group pages.
The operating temperature was changed on >>> APRIL 23, 1994 <<< to reduce the CTE problem and the number of hot pixels (see the Holtzman et. al., paper for details). This date represents a major change in the calibration of the WFPC2, since many instrument characteristics were affected by this temperature change (e.g., sensitivity, dark current, hot pixels, bias level, etc.).
1994 December 26 -76° C April 23 -88° C
OPUS pipeline software has been updated to fix a bug in the computation of statistics of raw (d0*) data as well as populate the FILTROT keyword (the rotation angle for partially rotated filters such as the ramps or polarizers).
2001 July 18
Alternate WFPC2 flatfields, with reduced noise, are available for manual recalibrations. Note that these are NOT in the pipeline as of Sep 2001.
2001 May 16
OTFR (On-The-Fly Reprocessing) is publically released. OTFR can be accessed via the HST Archive or STARVIEW. NOTE: uchcoord no longer needs to be run on OTFR data; for details as well as a comparison of OTFC and OTFR, see OTFR page.
A minor bug affecting only short INTFLATs has been fixed in the OPUS pipeline (pre-calwp2) software. Archival exposures will be fixed automatically by OTFR (On-The-Fly Reprocessing), planned for release in early 2001. The problem was that INTFLATs with non-integer exposure times have incorrect integer exposure times in the headers. More details and a full listing are given in this problem report.
Minor bug in calwp2 affects bias level calculation in AREA mode images only. Typically, the errors in the bias values are ~0.01 DN, though for a very small number of images (~20, those with large numbers of saturated pixels), the error can be ~1 DN or larger. Note however, that there are very few affected exposures: out of a total of nearly 90,000 WFPC2 images in the archive, only ~150 are AREA modes. All Cycle 7 and Cycle 8 GOs using AREA mode have been notified and CALWP2 will of course be updated as soon as possible. In the meantime, for assistance with any AREA mode images, please contact firstname.lastname@example.org.
OTFC (On-The-Fly Calibration) is publically released.
1999 Apr. - Jul.
Single chip data was processed as PC, regardless of the actual chip that was read out: generic conversion (the step preceding calwp2) was erroneously setting detector to 1 for all images with less then 4 groups. A software fix was installed on July 08, 1999. This mode is rarely used; no GO images were affected, only WFPC2 calibration data were corrupted (17 images total).
A new version of the OPUS software was installed in the pipeline Apr. 05, 1999, as part of switching the OPUS platform from VMS/VAX to ALPHA/VMS. The data processing steps remain the same however, there have been some minor keyword changes. The changes should be transparent to the majority of WFPC2 users, further details may be found in the software update summary.
Update to SYNPHOT contamination tables. Note however, that the pipeline does not automatically correct for contamination (e.g., the PHOTFLAM keyword). The correction must be applied manually, using either the information in Tables 2 and 4 of ISR98-03: WFPC2 Longterm Photometric Stability or via SYNPHOT (see section 4 of ISR 98-03).
STARVIEW and bestref now provide most recent GRAPHTAB and COMPTAB names, for use by calwp2 in computing the PHOTFLAM and other photometry-related keywords in the calibrated header files. (See also Note 2.)
1997 Aug. 15
A rare problem in the population of the PA_V3 keyword has been found, however, only 50 observations are affected. For more details, including a list of affected images and a fix for those images, please refer to our online PA_V3 Memo.
1997 May 16
SYNPHOT tables updated. With a few exceptions, changes to non-UV filter modes were relatively minor: generally ~1-2% or less. F785LP, F850LP, and F1042M were more: ~1-4%, ~1-6%, and ~2-15% for them, respectively, depending on the chip. The UV filters required somewhat larger changes to bring SYNPHOT into agreement with the observations, ranging from ~2% (e.g., F255W and F300W) to ~5% (F170W, F218W) to 10% or more (F160BW, F375N). Any filters not updated at this time will be checked with future observations (planned for summer of 1997) and updated if necessary. An ISR is expected out end of June. A table summarizing the photflam changes and new zero-points, as well as the README file (guide to WFPC2 SYNPHOT tables) and ASCII versions of the tables, can all be found in the Photometry and SYNPHOT Page. See also the May 1997 STAN and ISR97-10: WFPC2 Synphot Update and Note 2 below.
1997 Jan. 03
Calwp2 update (184.108.40.206): fix problem with processing of images containing less than 4 groups. Since Dec 1994 (220.127.116.11), observations that read out less than 4 chips AND whose first chip was not PC have been incorrectly calibrated as if all chips were PC. The problem has been identified, current GOs notified, the code fixed and installed. The problem did not affect 4-group images.
New flatfields for polarizers installed in pipeline. Filters include polarizers (i.e., polq, polqn33, polqn18, or polqp15) used with one of the following:
f300w, f336w, f390n, f410m, f439w, f547m, f555w, f606w, f656n, f675wFilenames are given in the WFPC2 Reffiles Memo.
1996 Aug. 15
Fix for WFPC2 images taken as a coordinated parallel with FOS (only). From Dec 95 to Aug 22,1996, WFPC2 images taken in parallel with the FOS were inadvertantly assigned FOS instead of WFPC pointing information. Note: these images have been fixed via reprocessing and rearchived as of May 1998; for a historical record of the problem, see online WFPC2 parallels fix memo.
1996 Aug. 09
New version of calwp2 (18.104.22.168; no change to data). Software now prints message to trailer file, header, and screen, if bias jump of greater than 0.09DN is found (jumps greater than 0.5DN have these messages as well as a WARNING printed to the PDQ files). For more details, please refer to the ISR 97-04: Properties of WFPC2 Bias Frames.
1996 ~Aug. 01
Darks are now generated in a new fashion; most importantly, more pixels are being flagged in the data quality files. Please refer to the short Summary of WFPC2 darks for details; an Instrument Science Report is being prepared. A Summary of WFPC2 Darks is also available.
SYNPHOT tables updated for FQCH4N* and FQUVN*. Affects only filter curves for 6193A and 8929A methane quad filters and the UV quad filters. Adjustments consisted of 3-15% decreases in filter throughputs. More details are available in the June 1996 STAN as well as the WFPC2 SYNPHOT Readme File.
Improved Cycle 4 flatfields installed in pipeline. Errors at edges now < 0.5% rms. Filters include all those redward of f300w:
f300w, f336w, f343n, f375n, f380w, f390n, f410m, f437n, f439w, f450w, f467m, f469n, f487n, f502n, f547m, f555w, f569w, f588n, f606w, f622w, f631n, f656n, f658n, f673n, f675w, f702w, f785lp, f791w, f814w, f850lp, f953n, f1042mFilenames are given in the WFPC2 Reffiles Memo.
1996 Apr. 18
WFPC2 headers from dithered data updated (dithered using scans, dither-line, or dither-box). Previously, due to the relatively small number of scans being performed, the pipeline provided only minimal support for WFPC2 dwell scans: only the initial dwell point position was used for the target pointing, so each exposure had the same CRVAL data, resulting in slightly incorrect metric results on all but the initial scan image. Note: the fix was only necessary for scan data; manually dithered images, e.g. using the POS-TARG option, had headers which properly reflected the pointing.
New flats for f300w, f375n, f450w, f555w, f606w, and f814w were installed in the pipeline. These are similar to the Cycle 4 flatfields but have been improved at the CCD edges: errors are now estimated to be < 1% rms. Filenames are given in the WFPC2 Reffiles Memo or will be found by Starview's 'bestref' screen.
Minor update to calwp2 (V.22.214.171.124): now prints out reference file names to go along with PEDIGREE/DESCRIP. Does not affect data; produces new informational messages during pipeline processing.
1995 Aug. 10
New maskfile (r0h) used in pipeline; columns above traps now flagged with "256" in c1h files. See ISR 95-03 for details.
WFPC2 related SYNPHOT tables updated. See Note 2, and the July 1995 STAN.
1995 March 24
There are 184 WFPC2 images, dating from Apr. '94 - Mar. '95, which have corrupted FILT* header keywords. Only those images taken with rotated filters (e.g., fqch4n, polqn33, polqp15, fquvn33, fr418n33, etc.) were affected. There were 2 sorts of problems: 1) when rotated filters were taken with a second (non-rotated) filter, the second name was erased from the header, and 2), when a rotated filter was used in isolation, OPUS would erroneously add some second filtername to the header. For more information, see the Rotated Filter Problem memo.
Observations taken 4/15/94 to 3/27/95, using WF3-FIX (only) may have incorrect pointing information in their image headers. In particular, errors on the order of ~2" may be present. Note that this only impacts 2 values in the header (although it can cause derived positions to be in error); there are no problems with the image data or with the telescope pointing. Note: these images have been fixed via reprocessing and rearchived as of May 1998; for a historical record of the problem, see online WF3-FIX Memo.
Minor update to 'calwp2' (V.126.96.36.199): now checks the PEDIGREE and DESCRIP keywords within reference files. Does not affect data; produces new informational messages during pipeline processing.
1994 May 04
Switch to separate values for even and odd columns. (Version 188.8.131.52 of 'calwp2'; note that new biasfiles are required with the new 'calwp2'. See Note 1.)
1994 Apr. 14
Fix installed for parallel WFPC2 data headers. Note that some archival data may still be incorrect. (See Note 3.)
1994 March 08-09
Prelaunch flats were used until this date. These were generally good to a few percent over most of the field, with a maximum deviation of 10 %. See Holtzman et al., paper for details.
1994 March 08
Overscan region used to determine the blev correction changed from [3:14,*] to [9:14,10:790]. (Version 184.108.40.206 of 'calwp2'. See Note 1.)
NOTE 1. Change to Overscan Region used for BLEVCORR
Until March 08, all 12 columns of the overscan region were used in the pipeline to determine the bias level. On March 08, a change was made to use only 6 columns, since the first few columns are too high or even saturated, having been affected by the exposure level. This results in the removal of too much bias level; a quick check of the DEZERO group keyword will indicate the amount removed, normal levels are ~300 to 400 DN.
There is also a very slight difference in bias level for the even and odd columns, typically about 0.1 DN difference although in WF4 it can be as high as 0.4 DN. Until May 04, 1994, a single value was used in the pipeline; separate values have been used in the pipeline since that time (these values are stored in the BIASEVEN and BIASODD group keywords in the science data headers). Note: the trailer (trl) file includes the calwp2 version that was used at the time of processing.
NOTE 2. New SYNPHOT Tables Available for recomputing PHOTFLAMs
Observers using the header photometry keywords (i.e., PHOTFLAM) from data processed in the pipeline before May 1997 should recompute these keywords, using either SYNPHOT directly or calwp2 (after updated SYNPHOT tables are retrieved and installed at their local site), or the SYNPHOT PHOTFLAMs and zeropoints table.
To identify epoch of SYNPHOT tables used for your observations: check the GRAPHTAB and COMPTAB keywords in the science data headers (c0h). The graphtab/comptab names used in the pipeline are based on the time of installation; for the May 1997 update, the GRAPHTAB and COMPTAB names are h5* or later (h=year since 1980, 5 is month, d is day. If the GRAPHTAB and COMPTAB names are h5* or later, the photometry keywords are already uptodate. If they predate May 1997, then the keywords should be recomputed. For more details about changes, installing files, and recomputing the keywords, refer to the Guide to WFPC2 Synphot Tables.
NOTE 3. Incorrect Pointing Information in some WFPC1 and WFPC-2 Parallel Data Headers;
Approximately 150 WFPC-1 and WFPC2 observations, taken in parallel with the FGS between Oct. 11, 1993 and Apr. 19, 1994, may have incorrect pointing information in their headers. Not all the data prior to the Apr 19 fix is affected, since workaround tools were developed after Mar 3 to allow PODPS operations personnel to fix the data but some observations were missed. The problem was related to the FGS observations, for which PODPS received the wrong roll orientation data. Some of these roll errors were significant and would affect the validity of the CD matrix. The images will be reprocessed and rearchived (expected completed by May 1996); if a list of affected observations is desired, please contact the STScI help desk at email@example.com.
Safings are usually not a major problem for WFPC2, which is placed in HOLD mode and does not warm up for most spacecraft safing situations. Even if the instrument should warm up and require a subsequent decontamination procedure, this is not an issue (as it was with WFPC 1) since there is no UV flood to preserve (although as with all decons, the UV calibration will change, see Section 1).
|HST entered inertial hold at 2006.107:03:10:00.5 after failing a bright earth/moon avoidance test. At the time of the event, ACS and WFPC2 were taking darks. The cause was determined to be related to a command scheduling logic problem between the FGS ITS (Internal Test Source) test activity and the standard PCS termination activity which must occur at the start of earth and moon occultation events. The problem prevented the scheduling of the PCS termination commanding which in turn allowed the FGS high voltage to remain turned on during the occultation event that followed the FGS ITS activities. The associated onboard Earth-Moon Bright Object Protection check failure condition properly initiated the immediate spacecraft transition to Inertial Hold. The current failure is strictly related to the FGS ITS test scheduling and has no implications for recovery of the science activities.|
|HST entered zero gyro safemode at 2004.188.16:28:51. Analysis indicates that this was NOT a result of problems with any of the gyros. The3 gyros are still working fine. The problem came from the loading of a table into the Flight Software that has data concerning the orientation and conversion factors for the gyros. The table was being reloaded to make very small adjustments in some of the parameters to improve the accuracy of the telescope slews. These parameters are also used by the Flight Software to automatically do sanity checks of the gyro data. The table got loaded with erroneous data, which resulted in the Flight Software test failing.|
|HST entered zero-gyro safe mode at 2003.119:14:21:42. Gyro three has stalled again. Attempts to restart gyro 3 have failed. Three successive attempts were made, but the gyro did not restart on any of them. While there will certainly be some follow up analysis of the engineering data from the stall and the restart attempts, gyro 3 is now considered failed. Gyro 1, one of the spares, was then started and it came on and has been operating properly. This event occurred during a scheduled WFPC2 decon, and resulted in a 32.7-hour decon (instead of a 6-hour one). After completion of the photometric monitor no adverse results were found.|
|HST entered Zero Gyro Software Sunpoint at 103/22:21:08 due to failure of System Momentum Test caused by an increase in Gyro 3 current GRG2_3MC, which reached a maximium value of 347.2 milliamps. Two Gyro Configuration Test subsequently failed at 22:24:46. Gyro 3 was successfully powered up on first attempt by Ops Request 16881-1 at 103/23:51:28. This is very similar to the occurrence on 11/28/2002.|
|HST experienced an electrical anomaly (short) on 2003.036. After examining data before and after the event no anomalous WFPC2 behaviour was detected.|
|HST went into Inertial Hold safemode at 11:59:18 UT (~08:00 local). The cause of the event was a missed command load; basically, HST ran out of things to do. There were no instrument or vehicle problems. This event occurred during a scheduled WFPC2 decon suite. The decon suite was repeated, and no negative effects were found.|
|HST went into zero gyro safe mode caused by a glitch in gyro 3. Gyro 3 has had several of these in the past year, but this is the first that has led to a safemode entry. HST was recovered from safemode 11/30 morning and resumed observations with no problems.|
|HST entered Zero Gyro Sun Point (ZGSP) mode when the gyro #5 motor currents exceeded the motor current safemode test limit. Science operations recommenced Apr 30.|
|HST entered software/sunpoint safemode at 06:38:06 UT (magnetic torquing bars were found to be out of limits). WFPC2 warmed to -40C; no decon was performed.|
|HST autonomously entered Software Sun Point (SWSP)
Safemode at 003/102417Z; WFPC2 DID NOT WARM UP.
The safemode was a result of failing the Solar Array Drive Electronics (SADE) Torque Test. Analyses indicate a SADE-2R failure early in a solar array slew. The anomaly had similar characteristics of a bit-flip event on day 090 of 1997. Upon entry to SWSP, SADE-1R was powered on and the arrays were successfully slewed to 90 degrees. Safemode recovery was successfully completed with the interception of the Health and Safety SMS at 004/0000Z and a test of SADE-2R was successfully performed at 004/0200Z. All SADE-2R telemetry was nominal as both positive and negative seven-degree slews were executed; flight software was configured for SADE-2R usage. Recovery of all science instruments will be accomplished by 004/1321Z and science observations are scheduled to begin execution at 004/1400Z.
|WFPC2 was intentionally safed (2000:306:22:35) as a precaution after additional shutter anomalies occurred with an apparent new type of behavior. The Anomaly Review Board recommendation for a software patch was implemented Nov. 08, 2000. WFPC2 was brought out of safe and into a nominal decon on Nov. 08, 2000 00:00 UT. For more information on the status of the shutter blade anomaly investigation, please see the shutter anomaly report.|
|After a series of shutter anomalies occurred, WFPC2 was intentionally safed (2000:276:11:47) while the problem was investigated. The conclusion was that the shutter was fine; instead, there may be a problem with a sensing circuit. An Anomaly Review Board has been formed to investigate further. WFPC2 was recovered from safe Oct. 05, 2000 at 21:16 UT; a normal decontamination procedure followed. For more information on the shutter blade anomaly, please see the shutter anomaly report.|
|HST entered PSEA (Pointing Safemode Electronics Assembly) safe mode at 17:06UT, due to a procedural error that occurred during a dump of the 486 memory. HST remained power positive and the aperture door open. Operations personnel transitioned to a Health and Safety SMS at 219/1000Z that remained in effect until 220/0000Z when a normal science load was resumed. WFPC2 was warm (-40C) for 2 days, which may change the contamination level. Standard star observations in the UV were taken Aug 7 (day 220), after WFPC2 was cooled down again; these will provide a check of the contamination level before WFPC2 enters the standard decontamination procedure on August 10. Details will be made available in the WWW standard star monitoring memo.|
|HST entered Zero Gyro safemode at 13:29 UT due to a gyro failure. Data acquistion was suspended until after the servicing mission in Dec 1999.|
|1999 May 07||HST enters inertial safemode at 1058Z, the Earth/Moon bright object detection test failed. WFPC2 did not warm up.|
|HST enters inertial hold safemode due to bad ephemeris load. WFPC2 did not warm up.|
|Safed due to gyro 4 failure (zero gyro hold, aperture door shut).|
|Safe (software sunpoint mode) due to solar array torque test failure during a routine SA maneuver for slew to target.|
|Warmed up to ~-50. HST in hardware sunpoint safemode caused by flight software bug. Decon executed Dec 18.|
|Did not warm up. Software sunpoint safe, due to reaction wheel problem, occurred during WFPC2 decon. The procedure, however, was unaffected: WFPC2 was cooled back down to -88° C with realtime commanding.|
|Did not warm up.|
|Did not warm up. Brief summary can be found in July 1996 STAN.|
|Did not warm up.|
|Did not warm up.|
|Did not warm up.|
The IRAF/STSDAS task "uchcoord" may be used to examine the plate scale of c0h files and/or update the CD matrices with a new plate scale; please see the online help for more details on this task.
The accuracy of the platescales, as derived from the fit, is 0.00002"/pixel for PC1 and about 0.00004 for the WF chips.
PC1 WF2 WF3 WF4 Nov. 1993 0.046 0.101 0.101 0.101 Apr. 11, 1994 0.0452 0.0991 0.0991 0.0993 Mar. 27, 1995 0.04553 0.09958 0.09956 0.09962 Apr. 14, 1996 0.045518 0.099548 0.099527 0.099587
NOTE 1: most images taken prior to April 11, 1994 will yield positional measurements ~3" in error on average, although it may be as much as 6". A software tool is being developed to aid in correcting the image headers for the pointing offset (and rotation) but until it is available, if you have any questions concerning the pointing of images taken in early 1994, please contact firstname.lastname@example.org. The rotation amounts to 0.8 deg for WF2 (smaller rotations for the other chips (0.28 deg in PC1, 0.46 deg in WF3, and 0.06 deg in WF4)). Since the STSDAS task METRIC uses WF2 as the reference chip, a 0.8 degree discrepancy will be present between data obtained before and after April 11, 1994, in addition to a shift.
NOTE 2: Images taken April 11-19, 1994 have positional information in the headers which does not correctly reflect the aperture updates. Affected GOs will be notified personally and an effort will be made to have the images reprocessed and re-archived. In the meantime, if you are working with images from the April 11-19, 1994 timeperiod, and have any questions concerning the pointing of images, please contact email@example.com.
table lists the distribution of WFPC2 modes in the HST Archive. Last update: February 2000.
- scale noise can be set separately for PC and WF - handles images with different exposure times - will create and/or update masks with cosmic rays flaggedFor more information about the changes and use of crrej, please refer to the online IRAF/STSDAS help. The previous version of crrej is still available as the STSDAS "mkdark". STSDAS 1.3.5 will be available from the STScI anonymous ftp site: ftp.stsci.edu (220.127.116.11) in the directory /software/stsdas/v1.3. Information on STSDAS 1.3.5 will also be posted to "What's New" on the STSDAS Home Page, and to the adass.iraf.announce newsgroup.
Note: this section 9 (Miscellaneous) has not been updated after Sep. 1998, if updates would be of use to you, please let us know via an Email to firstname.lastname@example.org.
The Keyword Exception Table is a listing of all observations in the HST Archive with incorrect keyword values, such as wrong filter number, shutter value, etc. When OTFC processes a dataset that is on this list, OTFC will automatically use the correct keyword value. This table is intended as a quick way for users to check if WFPC2 datasets have been added/removed from the list.
This table is a record of WFPC2 history in chronological order. See above for full details.
Last updated: Feb. 10, 2000.